# Imprecise Markov Semigroups and their Ergodicity

### v4

@@ 2. Imprecise Markov Semigroups
$\mathbf{P}\preceq^{\text{part}}_{\tilde{f}}\mathbf{P^{\prime}}\iff P_{t}\tilde{f}\leq P^{\prime}_{t}\tilde{f}\text{, }\forall t\geq 0,$
@@
but there might exist $\mathbf{P},\mathbf{P^{\prime}}\in\mathbf{Q}$ such that $\mathbf{P}\not\preceq^{\text{part}}_{\tilde{f}}\mathbf{P^{\prime}}$ and $\mathbf{P}\not\succeq^{\text{part}}_{\tilde{f}}\mathbf{P^{\prime}}$. Let us now introduce three set theoretic concepts (Wolk, 1958), that will allow us to properly define an Imprecise Markov Semigroup even when only a partial order can be defined.
Definition 2 (Up/Down-Directed Subsets of $\mathbf{Q}$).
@@
A subset $\mathbf{S}$ of $\mathbf{Q}$ is up-directed if for all $\mathbf{P},\mathbf{P^{\prime}}\in\mathbf{S}$, there exists $\mathbf{P^{\prime\prime}}\in\mathbf{S}$ such that $\mathbf{P}\preceq^{\text{part}}_{\tilde{f}}\mathbf{P^{\prime\prime}}$ and $\mathbf{P^{\prime}}\preceq^{\text{part}}_{\tilde{f}}\mathbf{P^{\prime\prime}}$. We say that $\mathbf{S}$ is down-directed if the opposite holds. That is, if for all $\mathbf{P},\mathbf{P^{\prime}}\in\mathbf{S}$, there exists $\mathbf{P^{\prime\prime}}\in\mathbf{S}$ such that $\mathbf{P^{\prime\prime}}\preceq^{\text{part}}_{\tilde{f}}\mathbf{P}$ and $\mathbf{P^{\prime\prime}}\preceq^{\text{part}}_{\tilde{f}}\mathbf{P^{\prime}}$.
Definition 3 (Dedekind-Closed Subsets of $\mathbf{Q}$).
@@
A subset $\mathbf{K}$ of $\mathbf{Q}$ is Dedekind-closed if whenever $\mathbf{S}$ is an up-directed subset of $\mathbf{K}$, and $\mathbf{P}$ is the $\preceq^{\text{part}}_{\tilde{f}}$-least upper bound of $\mathbf{S}$ (or $\mathbf{S}$ is a down-directed subset of $\mathbf{K}$, and $\mathbf{P}$ is the $\preceq^{\text{part}}_{\tilde{f}}$-greatest lower bound of $\mathbf{S}$), then $\mathbf{P}\in\mathbf{K}$.
@@
Definition 4 (Incomparability, Diversity, and $\preceq^{\text{part}}_{\tilde{f}}$-Width of $\mathbf{Q}$).
@@
Definitions 2, 3, and 4 are needed to introduce the concept of $\preceq^{\text{part}}_{\tilde{f}}$-compatible topology. Requiring that $\mathbf{Q}$ is compact in such a topology (together with another assumption) will ensure us that it possesses least and greatest elements (with respect to the partial order $\preceq^{\text{part}}_{\tilde{f}}$).
@@
A topology $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$ on $\mathbf{Q}$ is $\preceq^{\text{part}}_{\tilde{f}}$-compatible if
@@
  • every set of the form $\{\mathbf{P}\in\mathbf{Q}:\mathbf{P^{\prime}}\preceq^{\text{part}}_{\tilde{f}}\mathbf{P}\preceq^{\text{part}}_{\tilde{f}}\mathbf{P^{\prime\prime}}\}$ is $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-closed.
Lemma 6 (Topological Properties of $\mathbf{Q}$).
If $\mathbf{Q}$ has finite $\preceq^{\text{part}}_{\tilde{f}}$-width, then $\preceq^{\text{part}}_{\tilde{f}}$-compatible topology $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$ is unique. If in addition $\mathbf{Q}$ is $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-compact, then it possesses least and greatest elements (with respect to the partial order $\preceq^{\text{part}}_{\tilde{f}}$). We denote them by $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$, respectively.
@@
The first part of the lemma is a consequence of Wolk (1958, Theorem, page 528), which also shows that $(\mathbf{Q},\tau_{\preceq^{\text{part}}_{\tilde{f}}})$ is Hausdorff. The second part is an immediate consequence of requiring that $\mathbf{Q}$ is $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-compact. Indeed, from Definition 5, this entails that $\mathbf{Q}$ possesses least and greatest elements. ∎
@@
Let us add a discussion on why the conditions in Lemma 6 ensure us that the greatest and least elements for $\mathbf{Q}$ exist. Compactness with respect to the topology $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$ means that the limit points of the chains [^5] Recall that a chain in $\mathbf{Q}$ is a subset of $\mathbf{Q}$ that is totally ordered with respect to $\preceq^{\text{part}}_{\tilde{f}}$}. in the partially ordered set (poset) $\mathbf{Q}$ exist, and belong to the poset $\mathbf{Q}$ itself. In other words, the $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-compactness condition ensures that the poset $\mathbf{Q}$ does not “escape to infinity” in a topological sense.
@@
Finite $\preceq^{\text{part}}_{\tilde{f}}$-width restricts the degree of “spread” in the poset $\mathbf{Q}$, preventing it from having arbitrarily large levels of incomparability. In other words, it ensures us that the “pathological elements” of $\mathbf{Q}$ – i.e. the incomparable ones – are controllably few. This helps us closing the gap with respect to the total order $\preceq^{\text{tot}}_{\tilde{f}}$ case, for which no incomparable elements exist.
Let us add a simple example of a collection $\mathbf{Q}$ that satisfies (1) and has finite $\preceq^{\text{part}}_{\tilde{f}}$-width. Consider a finite collection $\mathbf{Q}=\{\mathbf{P}_{i}\}_{i=1}^{6}$ of six Markov semigroups $\mathbf{P}_{i}$ associated with as many Markov processes. Fix a functional $\tilde{f}\in B(E)$, and suppose that the partial order $\preceq^{\text{part}}_{\tilde{f}}$ can be defined, so that we can draw the following Hasse diagram
$\mathbf{P}_{6}$$\mathbf{P}_{2}$$\mathbf{P}_{3}$$\mathbf{P}_{4}$$\mathbf{P}_{5}$$\mathbf{P}_{1}$
Notice that, from top to bottom, a line segment represents the concept “is preferred to according to the partial order $\preceq^{\text{part}}_{\tilde{f}}$”, and if no such a segment connects two elements, this means that they are not comparable. So, for instance, $\mathbf{P}_{2}\preceq^{\text{part}}_{\tilde{f}}\mathbf{P}_{6}$ and $\mathbf{P}_{1}\preceq^{\text{part}}_{\tilde{f}}\mathbf{P}_{2}$, but $\mathbf{P}_{2}$ and $\mathbf{P}_{3}$ are incomparable. Then, the $\preceq^{\text{part}}_{\tilde{f}}$-width of $\mathbf{Q}$ is $4$, since the diverse set having the largest cardinality is $\mathbf{S}=\{\mathbf{P}_{2},\mathbf{P}_{3},\mathbf{P}_{4},\mathbf{P}_{5}\}$. In addition, we see that $\mathbf{Q}$ has a least element $\mathbf{P}_{1}$ and a greatest element $\mathbf{P}_{6}$; it is also trivially $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-compact since it is finite.
Together, $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-compactness and finite $\preceq^{\text{part}}_{\tilde{f}}$}-width ensure the existence of the greatest and least elements $\overline{\mathbf{P}}$ and $\underline{\mathbf{P}}$, respectively, in the poset $\mathbf{Q}$. We are now ready for the full definition of an Imprecise Markov Semigroup.
Definition 7 (Imprecise Markov Semigroup).
@@
Let $\mathbf{Q}$ be a subset of the space $({B(E)}^{\mathbb{R}_{+}})^{B(E)}$ that satisfies (1). Fix any $\tilde{f}\in B(E)$. If we can define the total order $\preceq^{\text{tot}}_{\tilde{f}}$ on $\mathbf{Q}$, then we say that $\mathbf{Q}$ is an Imprecise Markov Semigroup if it is $\tau_{\preceq^{\text{tot}}_{\tilde{f}}}$-compact.
@@
If instead we can only specify the partial order $\preceq^{\text{part}}_{\tilde{f}}$, then we say that $\mathbf{Q}$ is an Imprecise Markov Semigroup if it has finite $\preceq^{\text{part}}_{\tilde{f}}$-width, and is $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-compact.
@@
In both cases, we denote by $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$ its least and greatest elements, respectively.

### v5

@@ 2 Imprecise Markov Semigroups
$\mathbf{P}\preceq_{\tilde{f}}\mathbf{P}^{\prime}\quad\Longleftrightarrow\quad P_{t}\tilde{f}\leq P_{t}^{\prime}\tilde{f}\text{, }\forall t\geq 0.$
This preorder is antisymmetric on $\mathbf{Q}$ (hence a genuine partial order on the semigroups in $\mathbf{Q}$) if and only if the orbit map
$\mathbf{P}\mapsto(P_{t}\tilde{f})_{t\geq 0}\in B(E)^{\mathbb{R}_{+}}$
is injective on $\mathbf{Q}$, i.e. whenever $P_{t}\tilde{f}=P_{t}^{\prime}\tilde{f}$ for all $t\geq 0$, it follows that $\mathbf{P}=\mathbf{P}^{\prime}$ as semigroups. Equivalently, $\tilde{f}$ separates the elements of $\mathbf{Q}$ through their time evolution. In general, a single functional $\tilde{f}$ will not be separating, so assuming antisymmetry from $\tilde{f}$ can be a strong identifiability restriction on $\mathbf{Q}$. However, antisymmetry can hold in structured families (e.g. low-dimensional parametric classes) when $\tilde{f}$ is chosen so that the trajectories $t\mapsto P_{t}\tilde{f}$ uniquely identify the underlying semigroup within $\mathbf{Q}$. Identifiability of Markov kernel/semigroup can be guaranteed via a separating (measure-determining) class of test functions rather than on a single $\tilde{f}$ (Ethier and Kurtz, 1986; Blount and Kouritzin, 2010). In this paper we instead allow an arbitrary, application-driven choice of $\tilde{f}$ and handle potential non-antisymmetry by working with the quotient $\mathbf{Q}_{\tilde{f}}\coloneqq\mathbf{Q}/\sim_{\tilde{f}}$. In the special case where $\tilde{f}$ is separating on $\mathbf{Q}$, the equivalence relation $\sim_{\tilde{f}}$ is trivial and the quotient poset becomes unnecessary.
@@
Let us now introduce three set theoretic concepts, inspired by Wolk (1958), that will allow us to properly define a compact Imprecise Markov Semigroup even when only a partial order can be determined.
Definition 3 (Up/Down-Directed Subsets of $\mathbf{Q}_{\tilde{f}}$).
@@
A subset $\mathbf{S}$ of $\mathbf{Q}_{\tilde{f}}$ is up-directed if for all $\mathbf{P}],[\mathbf{P^{\prime}}]\in\mathbf{S}$, there exists $\mathbf{P^{\prime\prime}}]\in\mathbf{S}$ such that $\mathbf{P}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{P^{\prime\prime}}]$ and $\mathbf{P^{\prime}}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{P^{\prime\prime}}]$. We say that $\mathbf{S}$ is down-directed if the opposite holds. That is, if for all $\mathbf{P}],[\mathbf{P^{\prime}}]\in\mathbf{S}$, there exists $\mathbf{P^{\prime\prime}}]\in\mathbf{S}$ such that $\mathbf{P^{\prime\prime}}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{P}]$ and $\mathbf{P^{\prime\prime}}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{P^{\prime}}]$.
Definition 4 (Dedekind-Closed Subsets of $\mathbf{Q}_{\tilde{f}}$).
@@
A subset $\mathbf{K}$ of $\mathbf{Q}_{\tilde{f}}$ is Dedekind-closed if whenever $\mathbf{S}$ is an up-directed subset of $\mathbf{K}$, and $\mathbf{P}]$ is the $\preceq^{\text{part}}_{\tilde{f}}$-least upper bound of $\mathbf{S}$ (or $\mathbf{S}$ is a down-directed subset of $\mathbf{K}$, and $\mathbf{P}]$ is the $\preceq^{\text{part}}_{\tilde{f}}$-greatest lower bound of $\mathbf{S}$), then $\mathbf{P}]\in\mathbf{K}$.
@@
Definition 5 (Incomparability, Diversity, and $\preceq^{\text{part}}_{\tilde{f}}$-Width of $\mathbf{Q}_{\tilde{f}}$).
@@
Definitions 3, 4, and 5 are needed to introduce the concept of $\preceq^{\text{part}}_{\tilde{f}}$-compatible topology. Requiring that $\mathbf{Q}_{\tilde{f}}$ is compact in such a topology (together with other assumptions) will ensure us that it possesses least and greatest elements (with respect to the partial order $\preceq^{\text{part}}_{\tilde{f}}$).
@@
A topology $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$ on $\mathbf{Q}_{\tilde{f}}$ is $\preceq^{\text{part}}_{\tilde{f}}$-compatible if
@@
  • every set of the form $\{[\mathbf{P}]\in\mathbf{Q}_{\tilde{f}}:[\mathbf{P^{\prime}}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{P}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{P^{\prime\prime}}]\}$ is $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-closed.
The next remark explores the connections between the (generic) partial-order-compatible and the interval topologies.
Remark 2 (Interval Topology vs $\preceq$-Compatible Topologies).
Let $(X,\preceq)$ be a partially ordered set. Wolk (1958) considers two topologies naturally associated with the order.
(1) Interval topology. The interval topology $\mathcal{T}_{I}$ is the topology for which the family of order intervals
$a,b]\coloneqq\{x\in X:a\preceq x\preceq b\},\qquad a,b\in X,$
forms a closed subbasis. Equivalently, $\mathcal{T}_{I}$ is the coarsest topology on $X$ for which every interval $[a,b]$ is closed.
(2) $\preceq$-compatible topology. A topology $\mathcal{T}$ on $X$ is said to be $\preceq$-compatible if: (i) every $\mathcal{T}$-closed set is Dedekind-closed, and (ii) every interval $[a,b]$ is $\mathcal{T}$-closed.
Relationship. Condition (ii) implies that any $\preceq$-compatible topology must contain the interval topology,
$\mathcal{T}_{I}\subseteq\mathcal{T}.$
Moreover, if $\mathcal{T}_{D}$ denotes the topology whose closed sets are exactly the Dedekind-closed subsets of $X$, then condition (i) yields the upper bound
$\mathcal{T}\subseteq\mathcal{T}_{D}.$
Hence, $\preceq$-compatible topologies are precisely those lying between $\mathcal{T}_{I}$ and $\mathcal{T}_{D}$,
$\mathcal{T}_{I}\subseteq\mathcal{T}\subseteq\mathcal{T}_{D}.$
In particular, $\mathcal{T}_{I}$ is the minimal $\preceq$-compatible topology requirement (it enforces closed intervals only), whereas $\mathcal{T}_{D}$ is the maximal one (it enforces all Dedekind-closed sets to be closed). In special situations these two coincide (e.g. under finite $\preceq$-width, similarly to Lemma 7), yielding uniqueness of the $\preceq$-compatible topology.
The following lemma summarizes some useful topological (and order-theoretic) properties of $\mathbf{Q}$.
Lemma 7 (Topological Properties of $\mathbf{Q}_{\tilde{f}}$).
If $\mathbf{Q}_{\tilde{f}}$ has finite $\preceq^{\text{part}}_{\tilde{f}}$-width, then $\preceq^{\text{part}}_{\tilde{f}}$-compatible topology $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$ is unique. If in addition (a) $\mathbf{Q}_{\tilde{f}}$ is $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-compact, (b) $\preceq^{\text{part}}_{\tilde{f}}$ has closed graph, that is, the set
$\mathrm{Graph}(\preceq^{\text{part}}_{\tilde{f}})\coloneqq\{([\mathbf{P}],[\mathbf{P}^{\prime}])\in\mathbf{Q}_{\tilde{f}}\times\mathbf{Q}_{\tilde{f}}:[\mathbf{P}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{P}^{\prime}]\}$
is closed in the product topology $\tau_{\preceq^{\text{part}}_{\tilde{f}}}\times\tau_{\preceq^{\text{part}}_{\tilde{f}}}$, and (c) $\mathbf{Q}_{\tilde{f}}$ is both up- and down- directed, then $\mathbf{Q}_{\tilde{f}}$ admits $\preceq^{\text{part}}_{\tilde{f}}$-least and $\preceq^{\text{part}}_{\tilde{f}}$-greatest elements. We denote them by $\underline{[\mathbf{P}]}_{\tilde{f}}$ and $\overline{[\mathbf{P}]}_{\tilde{f}}$, respectively.
@@
The first part of the lemma is a consequence of Wolk (1958, Theorem, page 528), which also shows that $(\mathbf{Q}_{\tilde{f}},\tau_{\preceq^{\text{part}}_{\tilde{f}}})$ is Hausdorff.
For the second claim, assume that $\mathbf{Q}_{\tilde{f}}$ is $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-compact and that $\preceq^{\text{part}}_{\tilde{f}}$ has closed graph. Consider the family
$\mathcal{C}\coloneqq\{\downarrow[\mathbf{P}]:[\mathbf{P}]\in\mathbf{Q}_{\tilde{f}}\},\qquad\downarrow[\mathbf{P}]\coloneqq\{[\mathbf{R}]\in\mathbf{Q}_{\tilde{f}}:[\mathbf{R}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{P}]\}.$
Each $\downarrow[\mathbf{P}]$ is nonempty and closed: Indeed,
$\downarrow[\mathbf{P}]=\{[\mathbf{R}]\in\mathbf{Q}_{\tilde{f}}:([\mathbf{R}],[\mathbf{P}])\in\mathrm{Graph}(\preceq^{\text{part}}_{\tilde{f}})\},$
so it is the preimage of the closed set $\mathrm{Graph}(\preceq^{\text{part}}_{\tilde{f}})$ under the continuous map $\mathbf{R}]\mapsto([\mathbf{R}],[\mathbf{P}])$. Order $\mathcal{C}$ by reverse inclusion. Let $\mathcal{K}\subset\mathcal{C}$ be a chain for this order, i.e. a nested family of nonempty closed sets. By compactness, $\bigcap_{C\in\mathcal{K}}C\neq\emptyset$; Pick $\mathbf{P}]\in\bigcap_{C\in\mathcal{K}}C$. Fix $C\in\mathcal{K}$. Since $C=\downarrow[\mathbf{Q}_{0}]$ for some $\mathbf{Q}_{0}]\in\mathbf{Q}_{\tilde{f}}$ and $\mathbf{P}]\in C$, we have $\mathbf{P}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{Q}_{0}]$. If $\mathbf{R}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{P}]$, then by transitivity $\mathbf{R}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{Q}_{0}]$, hence $\mathbf{R}]\in C$. Thus $\downarrow[\mathbf{P}]\subseteq C$. Hence $\downarrow[\mathbf{P}]$ is an upper bound of $\mathcal{K}$ in the reverse-inclusion order. Therefore, by Zorn’s lemma, $\mathcal{C}$ has a maximal element $\downarrow[\mathbf{P}_{\ast}]$ (with respect to reverse inclusion). We claim that $\mathbf{P}_{\ast}]$ is $\preceq^{\text{part}}_{\tilde{f}}$-minimal in $\mathbf{Q}_{\tilde{f}}$. Indeed, if $\mathbf{R}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{P}_{\ast}]$, then $\downarrow[\mathbf{R}]\subseteq\downarrow[\mathbf{P}_{\ast}]$. By maximality of $\downarrow[\mathbf{P}_{\ast}]$ in $\mathcal{C}$ (reverse inclusion), we must have $\downarrow[\mathbf{R}]=\downarrow[\mathbf{P}_{\ast}]$, which implies $\mathbf{P}_{\ast}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{R}]$. By antisymmetry of $\preceq^{\text{part}}_{\tilde{f}}$ on $\mathbf{Q}_{\tilde{f}}$, $\mathbf{R}]=[\mathbf{P}_{\ast}]$. Hence $\mathbf{P}_{\ast}]$ is minimal. The existence of a maximal element follows analogously by applying the same argument to the closed principal upper sets $\uparrow[\mathbf{P}]\coloneqq\{[\mathbf{R}]\in\mathbf{Q}_{\tilde{f}}:[\mathbf{P}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{R}]\}$. Finally, assume $\mathbf{Q}_{\tilde{f}}$ is down-directed. Let $\mathbf{P}_{\ast}]$ be a $\preceq^{\text{part}}_{\tilde{f}}$-minimal element. For any $\mathbf{P}]\in\mathbf{Q}_{\tilde{f}}$, by down-directedness there exists $\mathbf{R}]\in\mathbf{Q}_{\tilde{f}}$ with $\mathbf{R}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{P}_{\ast}]$ and $\mathbf{R}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{P}]$. By minimality of $\mathbf{P}_{\ast}]$ we must have $\mathbf{R}]=[\mathbf{P}_{\ast}]$, hence $\mathbf{P}_{\ast}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{P}]$. Thus $\mathbf{P}_{\ast}]$ is the least element. The argument for a greatest element is analogous using up-directedness. ∎
@@
Let us add a discussion on why the conditions in Lemma 7 ensure that the greatest and least elements for $\mathbf{Q}_{\tilde{f}}$ exist. Compactness with respect to the topology $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$ can be read as a “no escape” condition: Whenever we approximate an extremal element by moving along a chain, the corresponding limit points – when they exist – remain inside $\mathbf{Q}_{\tilde{f}}$. In other words, $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-compactness ensures that $\mathbf{Q}_{\tilde{f}}$ does not “escape to infinity” in a topological sense: It prevents extremal candidates from slipping out of the model class.
@@
Finite $\preceq^{\text{part}}_{\tilde{f}}$-width restricts the degree of “spread” in the poset $\mathbf{Q}_{\tilde{f}}$, preventing it from having arbitrarily large levels of incomparability. In other words, it ensures that the “pathological elements” of $\mathbf{Q}_{\tilde{f}}$ – i.e. the incomparable ones – are controllably few. This helps bridge the gap with the totally ordered case $\preceq^{\text{tot}}_{\tilde{f}}$, where no incomparable elements exist and extremal elements are conceptually straightforward.
The closed-graph condition for $\preceq^{\text{part}}_{\tilde{f}}$ (on the quotient poset $\mathbf{Q}_{\tilde{f}}$) plays a complementary topological role. It says that comparability is stable under limits: If $\mathbf{P}_{n}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{P}_{n}^{\prime}]$ for all $n$ and $\mathbf{P}_{n}]\to[\mathbf{P}]$, $\mathbf{P}_{n}^{\prime}]\to[\mathbf{P}^{\prime}]$ in $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$, then necessarily $\mathbf{P}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{P}^{\prime}]$. In particular, each principal lower set $\downarrow[\mathbf{P}]\coloneqq\{[\mathbf{R}]\in\mathbf{Q}_{\tilde{f}}:[\mathbf{R}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{P}]\}$ and principal upper set $\uparrow[\mathbf{P}]\coloneqq\{[\mathbf{R}]\in\mathbf{Q}_{\tilde{f}}:[\mathbf{P}]\preceq^{\text{part}}_{\tilde{f}}[\mathbf{R}]\}$ is $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-closed. This prevents the order relation from “breaking” at the boundary of $\mathbf{Q}_{\tilde{f}}$ and is what allows compactness arguments (via intersections of closed sets) to produce genuine extremal elements inside $\mathbf{Q}_{\tilde{f}}$.
Finally, the up-directed and down-directed assumptions ensure that these extremal elements are not merely maximal/minimal in the partial order, but are in fact greatest/least. Down-directedness means that any two models admit a common lower bound in $\mathbf{Q}_{\tilde{f}}$; Thus, once a $\preceq^{\text{part}}_{\tilde{f}}$-minimal element $\underline{\mathbf{P}}]_{\tilde{f}}$ exists, it must lie below every other element (otherwise one could find a strict lower bound for it), hence it is the least element. The analogous statement holds for the greatest element under up-directedness. In short, compactness and closed graph yield existence of minimal/maximal elements, while two-sided directedness upgrades them to least/greatest, which is precisely what we need to identify the extremal elements within $\mathbf{Q}_{\tilde{f}}$.
We now give a simple example of a collection $\mathbf{Q}$ and a test function $\tilde{f}$ that satisfy (1) and the hypotheses of Lemma 7, without the additional complications arising from equivalence classes.
Example 8 (A Compact Family with Closed Order and Least/Greatest Elements).
Let $E=\{0,1\}$ with the discrete $\sigma$-algebra $\mathcal{F}=2^{E}$, and fix the test function $\tilde{f}:E\to\mathbb{R}$ given by $\tilde{f}(x)=x$. For each parameter $a\in[0,1]$, let $\mu_{a}$ be the Bernoulli(a)(a) law on $E$, i.e.
$\mu_{a}(\{1\})=a,\qquad\mu_{a}(\{0\})=1-a.$
Define the Markov operator $K_{a}$ by
$(K_{a}g)(x)\coloneqq\int_{E}g(y)\mu_{a}(\mathrm{d}y),\qquad g\in B(E),$
so that $K_{a}g$ is constant in $x$. Notice that $K_{a}$ is idempotent, $K_{a}^{2}=K_{a}$.
For $t\geq 0$, define
$P_{t}^{a}\coloneqq e^{-t}I+(1-e^{-t})K_{a}.$
Then $\mathbf{P}^{a}=(P_{t}^{a})_{t\geq 0}$ is a Markov semigroup. Indeed, using $K_{a}^{2}=K_{a}$ and $IK_{a}=K_{a}I=K_{a}$,
$P_{t}^{a}P_{s}^{a}=\bigl(e^{-t}I+(1-e^{-t})K_{a}\bigr)\bigl(e^{-s}I+(1-e^{-s})K_{a}\bigr)=e^{-(t+s)}I+\bigl(1-e^{-(t+s)}\bigr)K_{a}=P_{t+s}^{a}.$
Let
$\mathbf{Q}\coloneqq\{\mathbf{P}^{a}:a\in[0,1]\}.$
With respect to the order $\preceq^{\mathrm{part}}_{\tilde{f}}$ (i.e. $\mathbf{P}\preceq^{\mathrm{part}}_{\tilde{f}}\mathbf{P}^{\prime}\iff P_{t}\tilde{f}\leq P_{t}^{\prime}\tilde{f}$ for all $t\geq 0$), we have
$P_{t}^{a}\tilde{f}(x)=e^{-t}\tilde{f}(x)+(1-e^{-t})\mu_{a}(\tilde{f})=e^{-t}x+(1-e^{-t})a,$
and therefore
$\mathbf{P}^{a}\preceq^{\mathrm{part}}_{\tilde{f}}\mathbf{P}^{b}\quad\Longleftrightarrow\quad a\leq b.$
Hence $\mathbf{Q}$ is a chain (so it has finite width $1$).
Endow $\mathbf{Q}$ with the topology induced by the identification $a\mapsto\mathbf{P}^{a}$. Then $\mathbf{Q}$ is compact and Hausdorff (homeomorphic to $[0,1]$), and the order graph is closed since it corresponds to
$\{(a,b)\in[0,1]^{2}:a\leq b\},$
which is closed in $0,1]^{2}$. In particular, every order interval $\{\mathbf{P}\in\mathbf{Q}:\mathbf{P}^{a^{\prime}}\preceq^{\mathrm{part}}_{\tilde{f}}\mathbf{P}\preceq^{\mathrm{part}}_{\tilde{f}}\mathbf{P}^{a^{\prime\prime}}\}$ is closed.
Finally, $\mathbf{Q}$ admits a least and a greatest element:
$\underline{\mathbf{P}}_{\tilde{f}}=\mathbf{P}^{0},\qquad\overline{\mathbf{P}}_{\tilde{f}}=\mathbf{P}^{1}.$
We are now ready for the full definition of a compact Imprecise Markov Semigroup.
Definition 9 (Compact Imprecise Markov Semigroup).
@@
Let $\mathbf{Q}$ be an IMSG, that is, a subset of the space $({B(E)}^{\mathbb{R}_{+}})^{B(E)}$ that satisfies (1). Fix any $\tilde{f}\in B(E)$, and recall the quotient poset $\mathbf{Q}_{\tilde{f}}=\mathbf{Q}/\sim_{\tilde{f}}$ introduced above. If the induced order $\preceq^{\text{tot}}_{\tilde{f}}$ on $\mathbf{Q}_{\tilde{f}}$ is total, then we say that $\mathbf{Q}$ is a compact Imprecise Markov Semigroup if $\mathbf{Q}_{\tilde{f}}$ is $\tau_{\preceq^{\text{tot}}_{\tilde{f}}}$-compact.
@@
If instead the induced order $\preceq^{\text{part}}_{\tilde{f}}$ on $\mathbf{Q}_{\tilde{f}}$ is only partial, then we say that $\mathbf{Q}$ is a compact Imprecise Markov Semigroup if $\mathbf{Q}_{\tilde{f}}$ has finite $\preceq^{\text{part}}_{\tilde{f}}$-width, it is $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-compact, it is both up- and down-directed, and $\preceq^{\text{part}}_{\tilde{f}}$ has closed graph.
@@
In both cases, we denote by $\underline{[\mathbf{P}]}_{\tilde{f}}$ and $\overline{[\mathbf{P}]}_{\tilde{f}}$ the $\preceq_{\tilde{f}}-$least and $\preceq_{\tilde{f}}-$greatest elements of $\mathbf{Q}_{\tilde{f}}$, respectively.
Remark 3 (Notation: Quotient Poset, Representatives, and Shorthands).
Fix $\tilde{f}\in B(E)$. Whenever we invoke order-theoretic or topological notions associated with $\preceq_{\tilde{f}}$, we work on the quotient poset $\mathbf{Q}_{\tilde{f}}=\mathbf{Q}/\sim_{\tilde{f}}$ endowed with the induced antisymmetric order, and we use bracket notation to emphasise equivalence classes. In particular, $\underline{[\mathbf{P}]}_{\tilde{f}}$ and $\overline{[\mathbf{P}]}_{\tilde{f}}$ denote the $\preceq_{\tilde{f}}-$least and $\preceq_{\tilde{f}}-$greatest equivalence classes in $\mathbf{Q}_{\tilde{f}}$.
When we drop the brackets and write $\underline{\mathbf{P}}_{\tilde{f}}$ or $\overline{\mathbf{P}}_{\tilde{f}}$, we mean that we have fixed (once and for all in the relevant argument) a representative Markov semigroup in the corresponding class, i.e. $\underline{\mathbf{P}}_{\tilde{f}}\in\underline{[\mathbf{P}]}_{\tilde{f}}$ and $\overline{\mathbf{P}}_{\tilde{f}}\in\overline{[\mathbf{P}]}_{\tilde{f}}$. Any operator-level statement involving $\underline{\mathbf{P}}_{\tilde{f}}$ or $\overline{\mathbf{P}}_{\tilde{f}}$ (e.g. generators, invariant measures, curvature conditions) is to be understood as a statement about the chosen representatives. For notational simplicity, in the remainder of this section, and in Sections 3 and 4, we adopt the following abuse of notation, $\underline{\mathbf{P}}_{\tilde{f}}\equiv\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}_{\tilde{f}}\equiv\overline{\mathbf{P}}$, with the understanding that the dependence on the chosen function $\tilde{f}$ is implicit.